OpenStudy (jacky90):

Two cyclists start at the same point and travel in opposite directions. One cyclist travels 6 mi/h slower than the other. If the two cyclists are 114mi apart after 3 hours, what is the rate of each cyclist? Will medal and fan

OpenStudy (agent0smith):

Make the faster guy have a rate r. The slower is then r-6. Distance = rate * time, and time is 3 hours: 3r + 3(r-6) = 114

OpenStudy (dumbcow):

distance = rate*time time = 3 hrs rate for cyclist 1 = r , cyclist 2 = r-6  solve for r